OpenStudy (anonymous):

I have solved for the differential equation, I just don't know how to get the final answer.. -ycosy+siny = -sin(x^-1) +c initial condition: y(1/pi) = pi What do I plug in where and what is my final answer in exact form?

OpenStudy (anonymous):

Think of it this way: If you had the solution \(y=x+C\) (for simplicity) with the initial value \(y(0)=1\), you would then replace every \(y\) you see with \(1\) and every \(x\) with 0. The same can be done for your equation. If \(y\left(\dfrac{1}{\pi}\right)=\pi\), then \[-\pi\cos\pi+\sin\pi=-\sin\left(\frac{1}{\pi}\right)^{-1}+C\]

OpenStudy (anonymous):

what is -sin(1/pi) inverse equal to?

OpenStudy (anonymous):

I need an exact form

OpenStudy (anonymous):

so far I have -pi= sin(1/Pi) +c

OpenStudy (anonymous):

Well, firstly you have \[-\sin\bigg[\left(\frac{1}{\pi}\right)^{-1}\bigg]\] which means you have the reciprocal of \(\dfrac{1}{\pi}\), or just \(\pi\): \[-\sin\pi\] Now what's the sine of \(\pi\)? What's the negative of that?

OpenStudy (anonymous):

wait a minute something is up

OpenStudy (anonymous):

shouldn't it be sin (pi^-1)

OpenStudy (anonymous):

that becomes sin (1/pi)

OpenStudy (anonymous):

No, the initial condition is \(y\left(\dfrac{1}{\pi}\right)=\pi\), meaning \(x=\dfrac{1}{\pi}\) when \(y=\pi\). The \(x^{-1}\) in the general solution is equivalent to \(\dfrac{1}{x}\), so \[\frac{1}{1/\pi}=\pi\]

OpenStudy (anonymous):

so what would be my final answer for C, I just want to be sure I got it right.

OpenStudy (anonymous):

\[-\pi\cos\pi+\sin\pi=-\sin\left(\frac{1}{\pi}\right)^{-1}+C\] reduces to \[-\pi(-1)+0=-\sin\pi+C\] then \[\pi=0+C~~\iff~~C=\pi\]

OpenStudy (anonymous):

oh I got -1 but I C how you got C.

OpenStudy (anonymous):

Thank you best responce

OpenStudy (anonymous):

You're welcome!